In the most recent election 20 percent of all eligible college student voted if a random sample of10 student were surveyed
1-find the probability that exactly half of them voted in the election

Answer by rothauserc(4718)

We can use the binomial probability formula to solve this problem since either a student voted or not voted
:
Probability(P) (k events in n trials) = nCk * p^k * (1-p)^(n-k), where nCk = n! / (k! * (n-k)!)
:
p = 0.20, k=5, n=10
:
P = 10C5 * (0.20)^5 * (1-0.20)^(10-5) = 0.0264
: